Find the energy, the frequency and the momentum of an X-ray photon of wavelength 0.10 nm.

Grade:11

1 Answers

Deepak Patra
9 years ago
λ = 01 nm a) Energy = hc/ λ = 1242 ev.nm/0.1 nm = 12420 ev = 12.42 Kev = 12.4 kev. b) Frequency = C/ λ = 3 * 10^8/0.1 * 10^9 = 3 *10^8/10^-10 = 3 * 10^18Hz c) Momentum = E/C = 12.4 * 10^3 *1.6 * 10^-19/3* 10^^8 = 6.613 * 10^-24 kg-m/s = 6.62 * 10^-24 kg-m/s